Question 7:

Consider a European call option and a European put option on a non dividend-paying stock. The price of the stock is $100 and the strike price of both the call and the put is $103, set to expire in 1 year. Given that the price of the European call option is $10.57 and the risk-free rate is 5%, what is the price of the European put option via put-call parity?

Question 8:

Suppose a trader buys a call option with a strike price of $30 and a premium of $1.58. When the option was purchased (three months previous), the stock traded for $31/share. At expiration, the stock traded for $18/share. What is the traders net profit or loss, per share? (Type just the number to two decimal places in the response box, without commas, dollar signs or percent signs. Do not enter commas but use negative sign if necessary, so for example "-1,234" would be "-1234").

Question 9:

Suppose a trader buys a put option with a strike price of $50 and a premium of $4.67. When the option was purchased (three months previous), the stock traded for $49/share. At expiration, the stock traded for $47/share. What is the traders net profit or loss, per share? (Type just the number to two decimal places in the response box, without commas, dollar signs or percent signs. Do not enter commas but use negative sign if necessary, so for example "-12.34").

Question 10:

Suppose a trader buys a put option with a strike price of $50 and a premium of $1.77. When the option was purchased (three months previous), the stock traded for $49/share. At expiration, the stock traded for $57/share. What is the traders net profit or loss, per share? (Type just the number to two decimal places in the response box, without commas, dollar signs or percent signs. Do not enter commas but use negative sign if necessary, so for example "-1,234" would be "-1234").

Please answer all 4 questions.
Answer #1

Q7

As per Put call Parity, Stock price + Put premium = Call premium + Present value of strike

Present value of strike =Strike price/(1+r)n

=103/(1.05)1 i.e. 98.10

As per equation,

100+ Put premium=10.57 +98.10

Put premium = 8.67

Q8 Purchase call option wtith strike price of 30 at $1.58

At expiry day price of stock is $18

So pay off is 0 since price is less then strike price.

Profit/(Loss) = Pay off -Premium i.e. 0-1.58 or (1.58)

Q 9 Purchase put option wtith strike price of 50 at $4.67

At expiry day price of stock is $47

Pay off = (Strike price- Stock price)

= 50-47 i.e.$3

Profit/(Loss) = Pay off -Premium i.e. 3-4.67 or (1.67)

Q10 Purchase put option wtith strike price of 50 at $1.77

At expiry day price of stock is $57

Pay off is 0 since price is more then strike price on expiry day.

Profit/(Loss) = Pay off -Premium i.e. 0-1.77 or (1.77)